# Can Test-time Computation Mitigate Memorization Bias in Neural Symbolic Regression?

### v1

@@ B.1 TPSR
We detail the exact procedure for computing the reward in TPSR. As explained in Section 5, the reward is mainly calculated based on the generated expression’s numerical accuracy, with additional consideration given to its complexity. In TPSR, a hyperparameter $\lambda$ controls the balance between fitting accuracy and complexity. Given a set of $n$ data points $\left\{\mathbf{x}_{i},y_{i}\right\}_{i=1}^{n}$, and a candidate prediction $\tilde{f}$, the reward $r(\tilde{f}(\cdot)\mid\mathbf{x},\mathbf{y})$ is calculated as follows:
@@ B.2.2 Inference
$k_{\mathrm{rand}}$ Size of randomly sampled subtree set $E^{\prime}_{\mathrm{rand}}$ $9$
@@
$l_{\mathrm{max}}$ Maximum length of a subtree’s representation $9$
@@
[…]
@@
procedure NSR-gvs-Inference($\mathcal{D}$)
@@
Appendix C Details for Experiments and Implementation
@@
In this section, we describe the details for the experiments conducted in Section 4, 5, and D. We also provide details regarding our implementation and the computational resources used in our experiments.
@@
We provide the model with $100$ data points in all experiments. We selected the range of the data support as follows: for the AI Feynman dataset, we used the support defined by the dataset itself. For all other datasets, we sampled the support range using the same procedure as used when generating the training data. For error bars, we report the standard deviation across three different random seeds. For the method combining NSR-gvs and TPSR, however, we conducted experiments with only a single seed due to the long inference time. Our implementation for data generation, model training, and related components is based on the original implementation of NSRwH  [^1] https://github.com/SymposiumOrganization/ControllableNeuralSymbolicRegression. For the TPSR experiments, we used the official implementation provided by the authors  [^2] https://github.com/deep-symbolic-mathematics/TPSR. For both implementations, we used the version of the implementation that was available on May 15. 2025. We trained and tested the model on a single NVIDIA A100 GPU. Training requires approximately 24 hours either for $1000$ epochs on a dataset with $100,000$ expressions or for $10$ epochs on a dataset with $10$ million expressions. The time required to generate a single expression at test time is less than one minute when using only NeSymReS or NSRwH, approximately $3$ to $10$ minutes with TPSR or NSR-gvs, and around $2$ to $5$ hours when combining TPSR with NSR-gvs.
Appendix D Additional Experiments
D.1 Trade-off Between Performance and Computational Cost
The results in Section 5 show how the relationship between memorization bias and numerical accuracy differ between various test-time strategies. However, test-time strategies also differ in terms of the computational cost required to generate an expression. In this section, we aim to better understand each test-time strategy by analyzing the trade-off between performance and the computational cost of expression generation. We also varied the beam size during decoding for NeSymReS, TPSR, and NSR-gvs for a more comprehensive analysis. We tested under the controlled setting described in Section 4, using the not_included dataset as the test dataset.
To measure the computational cost, we followed the approach of Shojaee et al. [23] and used the number of candidate expressions generated by the model during the generation of a single equation. For example, this value corresponds to the beam size in NeSymReS, the number of total rollouts multiplied by beam size in TPSR, and the number of iteration loops multiplied by beam size in NSR-gvs.
We present the results in Figure 5. It can be observed that, unlike NeSymReS—where larger beam size yields only limited reduction in memorization bias—TPSR and NSR-gvs achieve notable reductions in memorization bias at comparable computational costs. However, in terms of numerical accuracy, simply increasing the beam size in NeSymReS yields better performance than using NSR-gvs or TPSR at a comparable computational cost. The results support the conclusion in Section 5 that the reduction of memorization bias is only weakly correlated with numerical accuracy.
D.2 Can NSRwH Also Mitigate Memorization Bias?
When researchers in fields of natural sciences or engineering model their experimental data, they often make use of prior knowledge. For example, scientists may anticipate a symmetry between variables or predict that a particular operator appears in the mathematical laws describing the data. NSRwH [3] is a method that enables incorporating such prior knowledge into the NeSymReS model. The types of prior knowledge provided to the model include the following:
While prior knowledge is required beforehand to use NSRwH, it is a method that provides the model with additional information during inference, similar to TPSR and NSR-gvs. In this section, we test whether NSRwH can mitigate memorization bias when prior knowledge is provided. We obtained a NSRwH model by finetuning the NeSymReS model that we trained in Section 4. We froze the numerical encoder of the NeSymReS model, attached a symbolical encoder, and finetuned the model for $250$ epochs. We used the same training dataset as in Section 4 consisting of $100,000$ expressions; however, during fine-tuning, prior knowledge was extracted from the ground-truth expressions and fed into the symbolic encoder. At test time, we evaluated the NSRwH model under settings where each type of prior knowledge is provided individually, as well as under a setting where all types of prior knowledge are provided simultaneously. We follow the default settings of NSRwH to determine the amount of prior knowledge provided during test-time, and we used the not_included dataset as the test dataset. We set the beam size to $5$ and compare the results with those of NeSymReS, which is also configured with a beam size of $5$.
Figure 6 shows the results for this experiment. While providing complexity, symmetry, or absent subtrees mitigates memorization bias only to a limited extent, providing appearing subtrees or providing all properties significantly mitigates memorization bias. However, we also observe that the numerical accuracy of NSRwH decreases when provided with appearing subtrees or with all properties. This indicates a limitation of NSRwH when dealing with data not included in the training set. The results also show that not all kinds of additional data are effective for mitigating memorization bias.
D.3 Do Novel Expressions Contribute to Improvements in Numerical Accuracy?
@@
In Section 5, we saw that providing additional information to the model during inference can lead to generation of novel expressions. However, we also demonstrated that mitigating memorization bias does not necessarily lead to better numerical accuracy. In this section, we analyzed how much the novel expressions generated under each test-time strategy (including NSRwH) contribute to improvements in numerical accuracy, and present the corresponding results in Table 5. “Novel” indicates that the generated expression does not appear in the training data, while “Not Novel” means it does. The values indicate the percentage of expressions that satisfy each condition.
@@
The results show that for test-time strategies that are capable of mitigating memorization bias (strategies shown in bold), a large proportion of generated novel expressions do not perform well in terms of numerical accuracy. Especially for TPSR, hardly any of the novel expressions exhibit high numerical accuracy. This indicates the difficulty of generating appropriate expressions from an expanded search space. However, for strategies using NSR-gvs, novel expressions contribute to high accuracy to some extent, showing that additional information can be beneficial for both mitigating memorization bias and improving numerical accuracy in some occasions.
Appendix E Theoretical Background and Proof
@@ E.1 Preliminary
E.1.1 Circuit Complexity Classes
@@
Whether $\mathsf{TC}^{0}$ is a proper subset of $\mathsf{NC}^{1}$ is an open question, but it is widely believed that this is the case. For a more detailed and comprehensive introduction, we recommend reference to [2].
@@
The class $\mathsf{AC}^{0}$ consists of Boolean circuits of constant depth and polynomial size whose gates have unbounded fan-in and are restricted to the basis $\{\mathsf{AND},\mathsf{OR},\mathsf{NOT}\}$. Intuitively, $\mathsf{AC}^{0}$ captures extremely shallow parallel computation.

### v2

@@ B.1 TPSR
Figure 5: Overview of TPSR. Expressions are generated by MCTS with guidance from a NSR model. During the expansion phase, the next token to expand is decided based on the NSR model’s logits, and during the simulation phase, the NSR model simulates the hypotheses instead of random simulation.
TPSR is a method where expressions are generated by MCTS. MCTS consists of $4$ key steps, and here we supplement details for the steps that were not covered in the main text. We also provide an overview in Figure 5.
First, during the selection phase, the policy-guided upper confidence bound (P-UCB) heuristic (Silver et al., 2018) is used, where the NSR model is used as the policy.
Then, as described in Section 6, the NSR model decides the token to expand during the expansion phase and simulates the hypotheses during the simulation phase, playing a central role in the generation process. After the simulation phase, the reward is mainly calculated based on the generated expression’s numerical accuracy, with additional consideration given to its complexity. In TPSR, a hyperparameter $\lambda$ controls the balance between fitting accuracy and complexity. Given a set of $n$ data points $\left\{\mathbf{x}_{i},y_{i}\right\}_{i=1}^{n}$, and a candidate prediction $\tilde{f}$, the reward $r(\tilde{f}(\cdot)\mid\mathbf{x},\mathbf{y})$ is calculated as follows:
@@
Finally, during the back-propagation phase, the scores for each node are updated by taking the maximum between the current score and the newly obtained reward value.
@@ B.2.2 Inference
$k_{\mathrm{rand}}$ Size of randomly sampled subtree set $E^{\prime}_{\mathrm{rand}}$ 99
@@
$l_{\mathrm{max}}$ Maximum length of a subtree’s representation 99
@@
The overall algorithm during inference-time is detailed in 1. For the hyperparameters that characterize the policy, we use the values shown in Table 4, which were tuned via Bayesian optimization on $5$ randomly generated expressions. The function $\lfloor\cdot\rfloor$ indicates the floor function, which rounds down the input to its nearest integer.
@@
end function
main NSR-gvs-Inference($\mathcal{D}$)
@@
end main
Appendix C Details for Experiments, Implementation, and use of LLMs
@@
In this section, we describe the details for the experiments conducted in Section 5, 6, and F. We also provide details regarding our implementation and the computational resources used in our experiments.
@@
We provide the model with $100$ data points in all experiments. We selected the range of the data support as follows: for the AI Feynman dataset, we used the support defined by the dataset itself. For all other datasets, we sampled the support range using the same procedure as used when generating the training data. For error bars, we report the standard deviation across three different random seeds. For the method combining NSR-gvs and TPSR, however, we conducted experiments with only a single seed due to the long inference time. Our implementation for data generation, model training, and related components is based on the original implementation of NSRwH  [^1] https://github.com/SymposiumOrganization/ControllableNeuralSymbolicRegression. For the transformer4sr and TPSR experiments, we used the official implementation provided by the authors  [^2] https://github.com/omron-sinicx/transformer4sr,  [^3] https://github.com/deep-symbolic-mathematics/TPSR. For both implementations, we used the version of the implementation that was available on May 15. 2025. We trained and tested the model on a single NVIDIA A100 GPU. Training requires approximately 24 hours either for $1000$ epochs on a dataset with $100,000$ expressions or for $10$ epochs on a dataset with $10$ million expressions. The time required to generate a single expression at test time is less than one minute when using only NeSymReS or NSRwH, approximately $3$ to $10$ minutes with TPSR or NSR-gvs, and around $2$ to $5$ hours when combining TPSR with NSR-gvs.
We used large language models (LLMs) to aid writing and coding, where we mainly used Gemini 2.5 Flash and GPT-5 to generate code and check on errors in writing.
Appendix D Discussion Concerning the Definition of Reproduction Bias
We first restate the definition of reproduction bias introduced in Section 5.1. This definition was introduced to discuss the novelty of expressions in the NeSymReS setting without considering constants.
Definition D.1 (Reproduction bias).
Given an output token sequence $\mathbf{s}$, let $e=\operatorname{seq}^{-1}(\mathbf{s})$ be the original expression that is represented by $\mathbf{s}$. If $\operatorname{strip}(e)\in E_{\operatorname{templ}}$, we say $\mathbf{s}$ is a reproduction of expressions seen during training.
Here, we define reproduction bias in the transformer4sr setting. NeSymReS and transformer4sr are different in the way expressions are stored in the training data; in NeSymReS, constant-free expressions $e_{\operatorname{templ}}$ are collected to form a set of training expressions $E_{\operatorname{templ}}$, whereas in transformer4sr, full expressions including constants, denoted by $e^{*}$, are stored, and the training dataset consists of the set $E^{*}$ of such expressions.
For transformer4sr, we measured expression novelty both with and without considering constants. If we consider the set
$E^{*}_{\operatorname{templ}}=\{e^{*}_{\operatorname{templ}}\;|\;e^{*}_{\operatorname{templ}}=\operatorname{strip(e^{*})}\}\quad\text{for all}\quad e^{*}\in E^{*}$
the definition of reproduction bias without considering constants would be the same as Definition D.1. When considering constants, the definition of reproduction bias would be slightly changed as follows:
Definition D.2 (Reproduction bias considering constants).
Given an output token sequence $\mathbf{s}$, let $e=\operatorname{seq}^{-1}(\mathbf{s})$ be the original expression that is represented by $\mathbf{s}$. If $e\in E^{*}$, we say $\mathbf{s}$ is a reproduction of expressions seen during training.
Next, we discuss the problem of structural equivalence and functional equivalence in the transformer4sr setting. While we have defined reproduction bias based on structural equivalence, one may argue that we should define and measure reproduction based on functional equivalence. For example, $x_{1}(x_{1}+x_{2})$ and $x_{1}^{2}+x_{1}x_{2}$ are different expressions from a structural perspective, but they are functionally equivalent. We stated in Remark 5.2 that for the NeSymReS setting, structural equivalence and functional equivalence can be treated the same due to the operator choices (there is no possibility that, for example, $x_{1}(x_{1}+x_{2})$ is generated when $x_{1}^{2}+x_{1}x_{2}$ is included in the training dataset). In the transformer4sr setting, this is not always the case, since we do not restrict the operators in this setting.
However, even if we were to redefine reproduction bias from the perspective of functional reproduction bias, the claims of Section 5.2 would remain unaffected. This is because every expression regarded as novel under the definition of functional reproduction bias is already also regarded as novel under the definition of structural reproduction bias. Since the claim of 5.2 is that the proportion of expressions classified as novel is small for naive NSR methods, adopting the definition of functional reproduction bias would only further reduce that proportion, without altering the direction of the conclusions.
Appendix E Additional Related Work
The GP framework is a traditional and widely used framework for solving symbolic regression. The GP algorithm is a method based on evolutionary computation; initially, several mathematical expressions are formed randomly, and subsequently the expressions are “evolved” by operations such as recombining two expressions, mutating an expression, and eliminating inappropriate expressions (Burlacu et al., 2020; Schmidt and Lipson, 2009; Virgolin et al., 2019; Cranmer, 2023).
An example of using brute-force algorithms for symbolic regression is AI Feynman (Udrescu and Tegmark, 2020; Udrescu et al., 2020). In AI Feynman, neural networks were used to identify properties such as symmetry and separability within given numerical data. These properties were then used to recursively simplify the problem, ultimately reducing it to a form amenable to brute-force solutions.
Petersen et al. (2019) proposed Deep Symbolic Regression (DSR), a method that uses reinforcement learning to tackle symbolic regression. In this approach, the authors used a recurrent neural network (RNN) to generate equations as token sequences, with the parameters that govern the selection of the token learned through reinforcement learning. Studies such as Symbolic Physics Learner (Sun et al., 2022) and Reinforcement Symbolic Regression Machine (Xu et al., 2024) also use reinforcement learning, where Monte Carlo Tree Search (MCTS) is applied to discover expressions.
Some studies combine several approaches for symbolic regression. For example, neural-guided genetic programming (Mundhenk et al., 2021) integrates DSR and genetic programming (GP), while the Unified DSR Framework (Landajuela et al., 2022) combines GP, AI Feynman, DSR, linear models, and NSR.
Appendix F Additional Experiments
F.1 Do Novel Expressions Contribute to Improvements in Numerical Accuracy?
@@
In Section 6, we saw that providing additional information to the model during inference can lead to generation of novel expressions. However, we also demonstrated that mitigating reproduction bias does not necessarily lead to better numerical accuracy. In this section, we analyzed how much the novel expressions generated under each test-time strategy (including NSRwH) contribute to improvements in numerical accuracy, and present the corresponding results in Table 5. “Novel” indicates that the generated expression does not appear in the training data, while “Not Novel” means it does. The values indicate the percentage of expressions that satisfy each condition.
@@
The results show that for test-time strategies that are capable of mitigating reproduction bias (strategies shown in bold), a large proportion of generated novel expressions do not perform well in terms of numerical accuracy. Especially for TPSR, hardly any of the novel expressions exhibit high numerical accuracy. This indicates the difficulty of generating appropriate expressions from an expanded search space. However, for strategies using NSR-gvs, novel expressions contribute to high accuracy to some extent, showing that additional information can be beneficial for both mitigating reproduction bias and improving numerical accuracy in some occasions.
F.2 Trade-off Between Performance and Computational Cost
The results in Section 6 show how the relationship between reproduction bias and numerical accuracy differ between various test-time strategies. However, test-time strategies also differ in terms of the computational cost required to generate an expression. In this section, we aim to better understand each test-time strategy by analyzing the trade-off between performance and the computational cost of expression generation. We also varied the beam size during decoding for NeSymReS, TPSR, and NSR-gvs for a more comprehensive analysis. We tested under the controlled setting described in Section 5, using the not_included dataset as the test dataset.
To measure the computational cost, we followed the approach of Shojaee et al. (2023) and used the number of candidate expressions generated by the model during the generation of a single equation. For example, this value corresponds to the beam size in NeSymReS, the number of total rollouts multiplied by beam size in TPSR, and the number of iteration loops multiplied by beam size in NSR-gvs.
We present the results in Figure 6. It can be observed that, unlike NeSymReS—where larger beam size yields only limited reduction in reproduction bias—TPSR and NSR-gvs achieve notable reductions in reproduction bias at comparable computational costs. However, in terms of numerical accuracy, simply increasing the beam size in NeSymReS yields better performance than using NSR-gvs or TPSR at a comparable computational cost. The results support the conclusion in Section 6 that the reduction of reproduction bias is only weakly correlated with numerical accuracy.
F.3 Can NSRwH Also Mitigate Reproduction Bias?
When researchers in fields of natural sciences or engineering model their experimental data, they often make use of prior knowledge. For example, scientists may anticipate a symmetry between variables or predict that a particular operator appears in the mathematical laws describing the data. NSRwH (Bendinelli et al., 2023) is a method that enables incorporating such prior knowledge into the NeSymReS model. The types of prior knowledge provided to the model include the following:
While prior knowledge is required beforehand to use NSRwH, it is a method that provides the model with additional information during inference, similar to TPSR and NSR-gvs. In this section, we test whether NSRwH can mitigate reproduction bias when prior knowledge is provided. We obtained a NSRwH model by finetuning the NeSymReS model that we trained in Section 6. We froze the numerical encoder of the NeSymReS model, attached a symbolical encoder, and finetuned the model for $250$ epochs. We used the same training dataset as in Section 5 consisting of $100,000$ expressions; however, during fine-tuning, prior knowledge was extracted from the ground-truth expressions and fed into the symbolic encoder. At test time, we evaluated the NSRwH model under settings where each type of prior knowledge is provided individually, as well as under a setting where all types of prior knowledge are provided simultaneously. We follow the default settings of NSRwH to determine the amount of prior knowledge provided during test-time, and we used the not_included dataset as the test dataset. We set the beam size to $5$ and compare the results with those of NeSymReS, which is also configured with a beam size of $5$.
Figure 7 shows the results for this experiment. While providing complexity, symmetry, or absent subtrees mitigates reproduction bias only to a limited extent, providing appearing subtrees or providing all properties significantly mitigates reproduction bias. However, we also observe that the numerical accuracy of NSRwH decreases when provided with appearing subtrees or with all properties. This indicates a limitation of NSRwH when dealing with data not included in the training set. The results also show that not all kinds of additional data are effective for mitigating reproduction bias.
F.4 Numerical Accuracy in Practical Settings
Figure 8: Comparison of test-time strategies on the AI-Feynman dataset. The general pattern remains consistent to the simplified setting, where large beam search and NSR-gvs+TPSR perform well, and NeSymReS+TPSR perform poorly.
Since the experiments in the main text were conducted under a relatively small training dataset with restricted operators, we additionally evaluated and compared the numerical performance of the test-time strategies under conditions that better reflect practical applications. A total of $10$ million expressions were used to construct the training dataset, employing all operators described in Section 3 without any restriction on operator types. We trained both a NeSymReS model and a prompt-augmented model on this dataset for $10$ epochs. For the test datasets, we prepared the following three sets:
  • AI-Feynman. This dataset consists of $91$ equations with up to five independent variables, extracted from the AIFeynman database (Udrescu and Tegmark, 2020). It is commonly used in various studies to assess the performance of symbolic regression methods.
  • only_five_variables_nc. This dataset consists of expressions containing exactly five independent variables, making it a challenging dataset. The “nc” designation indicates that the expressions do not include constants, which simplifies the problems slightly; however, it remains more difficult than the first dataset. The dataset was constructed by sampling expressions from $p_{\mathcal{E}}$, filtering for expressions that include exactly five variables, and finally deleting its constants. This dataset is derived from the study of NSRwH (Bendinelli et al., 2023), and we use the first $100$ expressions for evaluation.
  • black-box. We also evaluated on numerical data collected from the real world, whose ground-truth expressions do not exist. We extracted 35 expressions from the black-box dataset in SR-Bench (La Cava et al., 2021) whose number of independent variables are five or less. The data are often noisy and may be sampled from a range different from the numerical data that the models were trained on, making the task challenging for the test-time computation methods.
Figures 8 demonstrates how the different test-time strategies perform under AI-Feynman, only_five_variables_nc, and black-box datasets. TPSR relatively performs slightly better than in the controlled setting; however, the general pattern of numerical accuracy remains consistent. These results demonstrate that, NSR-gvs is able to improve performance in practical settings, including those with noisy data. However, it can also be seen that even in practical settings, test-time strategies that mitigate reproduction bias do not always result in better performance.
F.5 Varying the Number of Epochs in transformer4sr
Figure 9: Reproduction bias in transformer4sr with varying epochs. Throughout the training process, the majority of both expression trees (both including and excluding constants) are copied directly from the training data.
In Section 5, we tested whether reproduction bias occurs in the practical setting of transformer4sr, while varying the training dataset size. Here, we present the results of an analysis of reproduction bias as a function of the number of training epochs. In the experiments reported in Section 5, the number of training epochs was fixed to $100$ following the original paper; therefore, we evaluate reproduction bias by varying the number of epochs from $10$ to $100$ in increments of $10$. The results are shown in Figure 9. While reproduction bias is slightly lower at smaller numbers of epochs, it remains pronounced throughout the training process, indicating that reproduction bias cannot be attributed merely to underfitting or overfitting.
F.6 Further Results on the baseline dataset
As described in Section 5, the empirical results show that the baseline dataset is a much more easier dataset compared to the not_included dataset with naive inference. Since the majority of expressions in the baseline dataset are expressions that were included in the training dataset, by applying test-time strategies to this dataset, we can test whether the test-time strategies help in reproducing the expressions from the training data.
In Figure 10 and Figure 11, we present the results concerning the numerical accuracy for various test-time strategies. We also tested with NSRwH as well as the test-time strategies described in Section 6. The results show that only beam search with a large beam size significantly improves performance, aligning with the previous result in Section 6 that providing the model with additional information often prevents the model from reproducing expressions from the training data.
Figure 10: Numerical accuracy for various test-time strategies on the baseline (and not_included) dataset. The y-axis shows the proportion of expressions that exceeded the $R^{2}$ thresholds of $0.5$, $0.9$, $0.95$, $0.99$, $0.999$, $0.9999$ and $0.99999$, respectively. Only beam search with a large beam size significantly improves performance for the baseline dataset.
Figure 11: umerical accuracy for NSRwH on the baseline (and not_included) dataset. The y-axis shows the proportion of expressions that exceeded the $R^{2}$ thresholds of $0.5$, $0.9$, $0.95$, $0.99$, $0.999$, $0.9999$ and $0.99999$, respectively. Adding additional information significantly curves the performance in the baseline dataset.
Appendix G Theoretical Background and Proof
Figure 12: Overview of the theoretical analysis. When assuming $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$, the last-token prediction problem is $\mathsf{NC}^{1}$-hard (due to reduction from Boolean formula value problem) while Transformers can only solve problems in $\mathsf{TC}^{0}$.
@@ G.1 Preliminary
G.1.1 Circuit Complexity Classes
@@
Whether $\mathsf{TC}^{0}$ is a proper subset of $\mathsf{NC}^{1}$ is an open question, but it is widely believed that this is the case. For a more detailed and comprehensive introduction, we recommend reference to (Arora and Barak, 2009).
@@
The class $\mathsf{AC}^{0}$ consists of Boolean circuits of constant depth and polynomial size whose gates have unbounded fan‐in and are restricted to the basis $\{\mathsf{AND},\mathsf{OR},\mathsf{NOT}\}$. Intuitively, $\mathsf{AC}^{0}$ captures extremely shallow parallel computation.